OpenStudy (anonymous):

Please help. find an equation for a line parallel to the line 5x-y=-10

OpenStudy (anonymous):

Well first put it into the form y=mx+b So subtract 5x on both sides. -y=-5x-10 Divide by -1 and you get y=5x+10 Now that we know that... We know that parallels have the same slope and in our case that's 5.

OpenStudy (anonymous):

im sorry i forgot to put the point (0,0)this is where im confused

OpenStudy (anonymous):

Okay well your slope is still 5! All we have to do is solve for b by putting (0,0) into the equation I wrote, but instead of 10 we will put b y=5x+b Now let's plug in (0,0) 0=5(0)+b 0=0+b So b equals 0 So if we write the parallel equation we get y=5x

OpenStudy (anonymous):

ok thank you thats what i got but i thought i was wrong ;)

OpenStudy (anonymous):

Well nope you were right!

OpenStudy (anonymous):

thank you very much!!!

OpenStudy (anonymous):

No problem!